Joe traveled against the wind in a small plane for 3 hours. The return trip with the wind took 2.8 hours. Find the speed of the wind if the speed of the plane in still air is 180mph.
Answer by edjones(7311)

d=distance t=time s=speed w=wind speed
d=ts
going: d=3(180-w)=540-3w
returning: d=2.8(180+w)=504+2.8w
since the distances are equal: 540-3w=504+2.8w
add 3w and subtract 504 from each side. 5.8w=36
w=6.2 mph approx.
Ed